OpenStudy (anonymous):

help!!!Given a soda can with a volume of 21 and a diameter of 6, what is the volume of a cone that fits perfectly inside the soda can? (Hint: only enter numerals in the answer blank). (will give medal) Numerical Answers Expected!

OpenStudy (anonymous):

@croslali help plz

OpenStudy (anonymous):

sigh*

OpenStudy (anonymous):

this will my last question for the day k?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

the answer is 7

OpenStudy (anonymous):

\[volume of cylinder =\pi r ^{2}h\] \[Volume of cone=\frac{ 1 }{ 3 }\pi r ^{2}h=\frac{ 1 }{3 }*21=7\] Because r and h are same.

OpenStudy (anonymous):

@luka1998 follow the same method in the prvious quest ... just remeber one thing ... the volume occupied by a cone is 1/3 that of a cylinder of the same height and radius

OpenStudy (anonymous):

thx @croslali @surjithayer

OpenStudy (anonymous):

yw